The random variable X is best described by a normal distribution with mean of 30 and standard deviation of 6. Find the z-score that corresponds to th following X values. a. X = 21 b. X = 26 c. X = 35 d. X = 42 e. X = 50


1
Expert's answer
2022-03-21T15:41:07-0400

"Z=\frac{x-\mu}{\sigma}"

"\mu=30"

"\sigma=6"

"Z(21)=\frac{21-30}{6}=-1.5"

"Z(26)=\frac{26-30}{6}=-0.67"

"Z(35)=\frac{35-30}{6}=0.83"

"Z(42)=\frac{42-30}{6}=2"

"Z(50)=\frac{50-30}{6}=3.33"